The historical sales for a certain model of a single serve coffee maker at a specialty cookware store in units is: January, 22; February, 17; March, 18; April, 24; May, 16, and June, 15. Use single exponential smoothing with α = 0.2, and compute the exponential smoothing forecasts for February through June. Do not round intermediate calculations. Round your answers to two decimal places.

Month Observation Forecast
January 22 22
February 17
March 18
April 24
May 16
June 15
Answer #1

α = 0.2

1-α = 1-0.2 =0.8

Fn = (An-1*α )+(Fn-1*(1-α))

Month Observation Calculation Forecast
January 22 22 22
February 17 =22*0.2+22*0.8 22
March 18 =17*0.2+22*0.8 21
April 24 =18*0.2+21*0.8 20.40
May 16 =24*0.2+20.40*0.8 21.12
June 15 =16*0.2+21.12*0.8 20.10